# Speed of sound peak in isospin QCD: a natural prediction of the Medium Separation Scheme

### v1

@@ Abstract
We present predictions for the zero-temperature equation of state at finite isospin density using the Nambu–Jona-Lasinio (NJL) model within the Medium Separation Scheme (MSS) — a scheme that explicitly disentangles medium effects from the ultraviolet divergent vacuum terms. Recent lattice QCD results reveal a nonmonotonic speed of sound ($c_{s}^{2}$) as a function of isospin chemical potential ($\mu_{I}$), exhibiting explicit violation of the conformal bound $c_{s}^{2}=1/3$. These findings have attracted significant theoretical interest, as established models — including the NJL model — failed to anticipate this behavior prior to lattice simulations. Conventional NJL implementations yield unphysical artifacts, often attributed to regularization scale sensitivity stemming from nonrenormalizability. However, in this work, we demonstrate that the standard NJL framework combined with MSS quantitatively reproduces state-of-the-art lattice data for isospin QCD.
I Introduction
Over the years, significant efforts have been devoted to studying the phase structure of strongly interacting matter, as described by quantum chromodynamics (QCD). The current understanding of the QCD phase diagram at finite temperature ($T$) and baryon density ($n_{B}$) reflects a rich interplay between experimental and theoretical approaches. At low baryon density and increasing temperature, heavy-ion collisions and lattice QCD (LQCD) simulations indicate a smooth crossover from hadronic matter to the quark-gluon plasma Gyulassy:2004zy ; Borsanyi:2020fev .
@@ I Introduction
The peak in $c_{s}^{2}$ lying above the conformal limit seems to be a general feature of dense QCD, as suggested by finite baryon density phenomenology (astrophysical observations) Bedaque:2014sqa ; Fukushima:2025ujk , two-color QCD studies Iida:2024irv , and also effects of the color superconducting phase Gholami:2024ety . In the case of finite isospin density, it was first reported from LQCD simulations in Ref. Brandt:2022hwy , with the peak in $c_{s}^{2}$ located at $\mu_{I}/m_{\pi}\approx 1.6$, with a maximum value $c_{s}^{2}\approx 0.56$. This qualitative result was once again obtained from LQCD calculations in Refs. Abbott:2023coj ; Abbott:2024vhj , although the location of the peak was shifted to higher densities at $\mu_{I}/m_{\pi}\approx 2.5$, while maintaining a maximum of similar height as the previous case. Within the same study, a unified model description combined the results of chiral effective theory ($\chi$EFT) Adhikari:2019zaj , continuum-extrapolated LQCD and pQCD Kurkela:2009gj ; Fujimoto:2023mvc through a Gaussian process (GP). As such, we have an ideal scenario to assess the predictions of the isospin-asymmetric NJL model.
In this paper, we show that the existence of the speed of sound peak can can be simply explained in the context of the NJL effective model when making use of the Medium Separation Scheme (MSS). The MSS provides a way to fully disentangle medium contributions from divergent vacuum contributions such that only the latter are regularized. This prescription has proven reliable and often necessary for the correct description of various phenomenologies Farias:2005cr ; Farias:2016let ; Duarte:2018kfd ; Das:2019crc ; Lopes:2021tro ; Azeredo:2024sqc . We argue that the existence of the speed of sound peak is a natural prediction of the MSS, which in addition shows excellent agreement with LQCD data from Refs. Abbott:2023coj ; Abbott:2024vhj and recovers the conformal limit at high densities, as opposed to the more standard way of regularizing ultraviolet (UV) terms in NJL type of models [^1] We call the usual form of regularizing UV divergent terms in the literature of Traditional Regularization Scheme (TRS)..
The remainder of this paper is structured as follows. In Sec. II, we briefly recapitulate the formalism of the NJL model with an isospin imbalance. The results obtained from the TRS and MSS procedures are presented and discussed in Sec. III. In Sec. IV, we synthesize our findings and emphasize the importance of MSS in NJL-type models to properly describe medium-dependent physical quantities.
II NJL Model within the Medium separation scheme
The NJL model with an isospin asymmetry has been extensively discussed in the literature Toublan:2003tt ; Frank:2003ve ; Barducci:2004tt ; He:2005sp ; He:2005nk ; He:2006tn ; Ebert:2005wr ; Ebert:2005cs ; Andersen:2007qv ; Sun:2007fc ; Abuki:2008wm ; Mu:2010zz ; Xia:2013caa ; Ebert:2016hkd ; Khunjua:2017khh ; Khunjua:2018sro ; Khunjua:2018jmn ; Khunjua:2019lbv ; Khunjua:2019ini ; Avancini:2019ego ; Lu:2019diy ; Khunjua:2020xws ; Lopes:2021tro ; Khunjua:2021oxf ; Liu:2021gsi ; Ayala:2023mms ; Carlomagno:2024xmi ; Liu:2023uxm . For further details on the medium separation scheme, see Refs. Duarte:2018kfd ; Lopes:2021tro , which have considered both a two-flavor and three-flavor analysis.
@@ II NJL Model within the Medium separation scheme
[…]
@@
The thermodynamic potential, obtained from the partition function as $\Omega=(\beta V)^{-1}\ln Z$, contains momentum integrals that are UV divergent and must be regularized. Given the nonrenormalizability of the NJL model, we introduce a cutoff $\Lambda$ that serves as an energy scale. The most common procedure in the literature (TRS), consists of simply introducing the UV cutoff $\Lambda$ in the divergent momentum integral. However, it is important to observe that the dispersion relations of the light quarks have the form:
@@
i.e, they are explicit functions of the isospin chemical potential. In Eq. (4), $E_{k}=\sqrt{k^{2}+M_{l}^{2}}$ and $M_{l}$ is the effective quark mass. In TRS, these terms are integrated up to $k=\Lambda$, and important medium effects are lost. In the next section, it will be evident that this approach introduces artifacts to the model, and the results cannot be trusted, especially in the high density regime Farias:2005cr ; Farias:2016let ; Duarte:2018kfd ; Das:2019crc ; Lopes:2021tro ; Azeredo:2024sqc . The MSS consists of manipulating the expressions for the gap equations  [^2] The gap equations are obtained by extremizing the thermodynamic potential, $\frac{\partial\Omega}{\partial M}=\frac{\partial\Omega}{\partial\Delta}=0$. All expressions for gap equations and thermodynamic potential at finite $\mu_{I}$ for $SU(2)$ and $SU(3)$ cases can be obtained for both TRS and MSS (see, e.g., Refs. Lopes:2021tro ; Avancini:2019ego for details. and thermodynamic potential by adding and subtracting convenient terms, and then rearranging the expressions separating divergences in terms of vacuum quantities, namely the vacuum effective quark mass $M_{0}=M(T=\mu_{I}=0)$. Such divergent integrals in the scale $M_{0}$ can then be regularized with the momentum cutoff, while the remaining medium terms are finite and can be integrated without restrictions. This procedure is crucial for a proper description of the thermodynamic quantities of interest, such as the pressure $P$, the isospin density $n_{I}$, the energy density $\epsilon$, and the speed of sound $c_{s}^{2}$, which are defined, respectively, as
@@
In the next section, we present the numerical results for each of these quantities.
III Numerical Results
To compare our model predictions with the results from LQCD, in the following we consider two representative sets of parameters: set I corresponds to a pion mass $m_{\pi}=139.57$ MeV, utilized in the model of Ref. Abbott:2024vhj , while set II corresponds to $m_{\pi}=169$ MeV to more closely represent the LQCD ensembles in Ref. Abbott:2023coj . The $SU(2)$ model takes as additional inputs the pion decay constant $f_{\pi}$ and the quark condensate $\left\langle\bar{u}u\right\rangle=\left\langle\bar{d}d\right\rangle$. The $SU(3)$ parametrization considers in addition to $f_{\pi}$ and $m_{\pi}$ also the kaon mass $m_{K}$ and the eta-meson mass $m_{\eta}$, while fixing the value of the current quark mass $m_{l}$. The obtained parameters are shown in Tabs. 1 and 2, for the $SU(2)$ and $SU(3)$ cases respectively.
@@ III Numerical Results
In Fig. 1, we show the behavior of the speed of sound as a function of the isospin chemical potential for both regularization prescriptions in the NJL model. Our results are compared with those obtained from the LQCD ensembles of Refs. Abbott:2023coj ; Abbott:2024vhj . To more closely match the value of $m_{\pi}$ used in the lattice, we have adopted the parameter set II in this case. Evidently, the MSS results show excellent agreement with the LQCD data, lying almost entirely within the error bands, in addition to describing a peak of similar height and even converging to the conformal limit at large $\mu_{I}$. The TRS description, on the other hand, is not able to describe the overall behavior because a finite value cannot be assigned to the speed of sound at large isospin chemical potentials. This behavior can be related to the pion condensate, which vanishes at a high $\mu_{I}$ in the TRS, while remaining nonvanishing within the MSS prescription, as shown in Fig. 2. This represents a very strong indication that the nonvanishing $\Delta$ at high values of $\mu_{I}$ obtained with the MSS is crucial to obtain the correct description in this regime.
@@
However, it should be noted that the reference LQCD data per ensemble may be subjected to finite lattice spacing effects and uncontrolled systematics at the low and high $\mu_{I}$ limits. For this reason, Ref. Abbott:2024vhj has considered a Bayesian model mixing framework, combining the results from pQCD, LQCD and $\chi$EFT through a Gaussian process in what is called the GP-model. A comparison between the NJL description and the GP-model is presented in Fig. 3. Since the corresponding pion mass is $m_{\pi}=139.57$ MeV in this case, we utilize the parameter set I. Once more, we see that only the MSS prescription can reproduce the qualitative behavior, although this time the speed of sound goes slightly above the reference data near the peak, and decreases slightly faster at intermediate isospin chemical potentials before converging to the conformal limit. Interestingly enough, the QMM results from Ref. Andersen:2025ezj (QMM A) and Ref. Brandt:2025tkg (QMM B) show the same overshoot and undershoot profile, and are in excellent agreement with the NJL results within the MSS. The fact that results obtained in a renormalizable model in the form of the QMM can be replicated by the MSS is another strong argument in favor of this regularization prescription. Also notable is the similarity between the $SU(2)$ and $SU(3)$ model descriptions, indicating that the strange quark plays a small role in this system. This can be anticipated, as the isospin imbalance has origin in the light quark sector. In regard to the TRS results, we once again cannot verify the existence of the peak or the convergence to the conformal limit. This is again a consequence of the vanishing of the pion condensate at high values of $\mu_{I}$, as seen in the previous case for the TRS. Furthermore, in the MSS, the height of the peak respects the upper bound $c_{s}^{2}=0.781$ of nuclear matter given in Ref. Hippert:2024hum , which was obtained through the calculation of hydrodynamic transport coefficients from first principles.
@@
IV Conclusions
@@
In this paper, we investigated the importance of the regularization scheme for the description of isospin imbalanced matter in the context of the NJL model. Of particular interest is the existence of a peak in the speed of sound, recently observed by lattice QCD simulations Brandt:2022hwy ; Abbott:2023coj ; Abbott:2024vhj . We worked within the MSS, which disentangles medium contributions from the vacuum such that only the latter is regularized. The results were compared with those obtained by the usual TRS procedure, which is typically used in NJL-type models in the literature. Throughout the analysis, it was made clear that the speed of sound peak is a natural prediction of the MSS. Of course, one must keep in mind that the NJL model has limited applicability at high energy scales as it is nonrenormalizable. Consequently, results with an isospin chemical potential greater than the regularization scale $\Lambda$ must be taken with great care. In fact, the TRS results are problematic at high $\mu_{I}$ and do not recover the conformal limit for the speed of sound. However, remarkably, despite the shortcomings expected from an effective model, the MSS results asymptotically approach $c_{s}^{2}=1/3$ as $\mu_{I}$ increases. This feature may be related to the behavior of the pion condensate, which is nonvanishing in this regime, in contrast to the prediction from the TRS prescription. Although concerned with superconducting systems, the Renormalization Group consistent regularization proposed in Ref. Gholami:2024diy finds a nonvanishing diquark condensate at large chemical potential, which is not the case with a conventional regularization. In this sense, the matching behaviors seem to signal that the MSS plays an important role in minimizing the dependence of physical quantities on the regularization scale. For this reason, we believe that a comparison with the RG-consistent method at finite isospin density would be interesting and could potentially help to further validate the applicability of the medium separation scheme also at the high density regime.
Acknowledgements.
@@
This work was partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), grants Nos. 312032/2023-4 (R.L.S.F.), 141270-2023-3 (B.S.L.) and 307286/2021-5 (R.O.R.); Fundação de Amparo à Pesquisa do Estado do Rio Grande do Sul (FAPERGS), Grants Nos. 24/2551-0001285-0 and 23/2551-0000791-6 (R.L.S.F.); Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021 (R.O.R.). The work is also part of the project Instituto Nacional de Ciência e Tecnologia—Física Nuclear e Aplicações (INCT—FNA), Grant No. 464898/2014-5 and supported by the Serrapilheira Institute (Grant No. Serra-2211-42230).

### v2

@@ Abstract
We present predictions for the zero-temperature equation of state at finite isospin density using the Nambu-Jona-Lasinio (NJL) model within the medium separation scheme (MSS)—a scheme that explicitly disentangles medium effects from the ultraviolet divergent vacuum terms. Recent lattice QCD results reveal a nonmonotonic speed of sound ($c_{s}^{2}$) as a function of isospin chemical potential ($\mu_{I}$), exhibiting explicit violation of the conformal bound $c_{s}^{2}=1/3$. These findings have attracted significant theoretical interest, as established models—including the NJL model—failed to anticipate this behavior prior to lattice simulations. Conventional NJL implementations yield unphysical artifacts, often attributed to regularization scale sensitivity stemming from nonrenormalizability. However, in this work, we demonstrate that the standard NJL framework combined with MSS quantitatively reproduces state-of-the-art lattice data for isospin QCD.
Introduction. Over the years, significant efforts have been devoted to studying the phase structure of strongly interacting matter, as described by quantum chromodynamics (QCD). The current understanding of the QCD phase diagram at finite temperature ($T$) and baryon density ($n_{B}$) reflects a rich interplay between experimental and theoretical approaches. At low baryon density and increasing temperature, heavy-ion collisions and lattice QCD (LQCD) simulations indicate a smooth crossover from hadronic matter to the quark-gluon plasma Gyulassy:2004zy ; Borsanyi:2020fev .
@@
[…]
In this paper, we show that the existence of the speed of sound peak can be simply explained in the context of the NJL effective model when making use of the medium separation scheme (MSS). The MSS provides a way to fully disentangle medium contributions from divergent vacuum contributions such that only the latter are regularized. This prescription has proven reliable and often necessary for the correct description of various phenomenologies Farias:2005cr ; Farias:2016let ; Duarte:2018kfd ; Das:2019crc ; Lopes:2021tro ; Azeredo:2024sqc . We argue that the existence of the speed of sound peak is a natural prediction of the MSS, which in addition shows excellent agreement with LQCD data from Refs. Abbott:2023coj ; Abbott:2024vhj and recovers the conformal limit at high densities, as opposed to the more standard way of regularizing ultraviolet (UV) terms in NJL type of models [^1] We call the usual form of regularizing UV divergent terms in the literature of traditional regularization scheme (TRS)..
The remainder of this paper is structured as follows. We first briefly recapitulate the formalism of the NJL model with an isospin imbalance. The results obtained from the TRS and MSS procedures are presented and discussed in sequence. Finally, we synthesize our findings and emphasize the importance of MSS in NJL-type models to properly describe medium-dependent physical quantities.
NJL Model within the Medium Separation Scheme. The NJL model with an isospin asymmetry has been extensively discussed in the literature Toublan:2003tt ; Frank:2003ve ; Barducci:2004tt ; He:2005sp ; He:2005nk ; He:2006tn ; Ebert:2005wr ; Ebert:2005cs ; Andersen:2007qv ; Sun:2007fc ; Abuki:2008wm ; Mu:2010zz ; Xia:2013caa ; Ebert:2016hkd ; Khunjua:2017khh ; Khunjua:2018sro ; Khunjua:2018jmn ; Khunjua:2019lbv ; Khunjua:2019ini ; Avancini:2019ego ; Lu:2019diy ; Khunjua:2020xws ; Lopes:2021tro ; Khunjua:2021oxf ; Liu:2021gsi ; Ayala:2023mms ; Carlomagno:2024xmi ; Liu:2023uxm . For further details on the medium separation scheme, see Refs. Duarte:2018kfd ; Lopes:2021tro , which have considered both a two-flavor and three-flavor analysis.
@@
$Z=\int[d\bar{\psi}][d\psi]\exp\left[-\int_{0}^{\beta}\!\!d\tau\!\int\!d^{3}x\left(\mathcal{L}+\hat{\mu}\bar{\psi}\gamma^{0}\psi\right)\right]\,$, (1)
@@
The thermodynamic potential, obtained from the partition function as $\Omega=(\beta V)^{-1}\ln Z$, contains momentum integrals that are UV divergent and must be regularized. Given the nonrenormalizability of the NJL model, we introduce a cutoff $\Lambda$ that serves as an energy scale. The most common procedure in the literature (TRS), consists of simply introducing the UV cutoff $\Lambda$ in the divergent momentum integral. However, it is important to observe that the dispersion relations of the light quarks have the form
@@
i.e., they are explicit functions of the isospin chemical potential. In Eq. (4), $E_{k}=\sqrt{k^{2}+M_{l}^{2}}$ and $M_{l}$ is the effective quark mass. In TRS, these terms are integrated up to $k=\Lambda$, and important medium effects are lost. In the next section, it will be evident that this approach introduces artifacts to the model, and the results cannot be trusted, especially in the high density regime Farias:2005cr ; Farias:2016let ; Duarte:2018kfd ; Das:2019crc ; Lopes:2021tro ; Azeredo:2024sqc . The MSS consists of manipulating the expressions for the gap equations  [^2] The gap equations are obtained by extremizing the thermodynamic potential, $\frac{\partial\Omega}{\partial M}=\frac{\partial\Omega}{\partial\Delta}=0$. All expressions for gap equations and thermodynamic potential at finite $\mu_{I}$ for $SU(2)$ and $SU(3)$ cases are shown in details in Refs. Lopes:2021tro ; Avancini:2019ego for both TRS and MSS cases. and thermodynamic potential by adding and subtracting convenient terms, and then rearranging the expressions separating divergences in terms of vacuum quantities, namely the vacuum effective quark mass $M_{0}=M(T=\mu_{I}=0)$. Such divergent integrals in the scale $M_{0}$ can then be regularized with the momentum cutoff, while the remaining medium terms are finite and can be integrated without restrictions.
As we will see, regardless of the form of regularization employed, the MSS is crucial for a proper description of the thermodynamic quantities of interest, such as the pressure $P$, the isospin density $n_{I}$, the energy density $\epsilon$, and the speed of sound $c_{s}^{2}$, which are defined, respectively, as
@@
In the next section, we present the numerical results for each of these quantities. It should be emphasized that an improper regularization of medium effects results in an incorrect behavior of thermodynamic quantities and may lead to nonphysical outcomes. A detailed discussion of causality violation and related issues arising from such improper treatments can be found in Ref. Pasqualotto:2023hho .
Numerical Results. To compare our model predictions with the results from LQCD, in the following, we consider two representative sets of parameters: set I corresponds to a pion mass $m_{\pi}=139.57$ MeV, used in the model of Ref. Abbott:2024vhj , while set II corresponds to $m_{\pi}=169$ MeV to more closely represent the LQCD ensembles in Ref. Abbott:2023coj . The $SU(2)$ model takes as additional inputs the pion decay constant $f_{\pi}$ and the quark condensate $\left\langle\bar{u}u\right\rangle=\left\langle\bar{d}d\right\rangle$. The $SU(3)$ parametrization considers in addition to $f_{\pi}$ and $m_{\pi}$ also the kaon mass $m_{K}$ and the eta-meson mass $m_{\eta}$, while fixing the value of the current quark mass $m_{l}$. The parameters obtained with the sharp cutoff method are shown in Tables 1 and 2, for the $SU(2)$ and $SU(3)$ cases respectively.
@@
In Fig. 1, we show the behavior of the speed of sound as a function of the isospin chemical potential for both regularization prescriptions in the NJL model. Our results are compared with those obtained from the LQCD ensembles of Refs. Abbott:2023coj ; Abbott:2024vhj . To more closely match the value of $m_{\pi}$ used in the lattice, we have adopted the parameter set II in this case. Evidently, the MSS results show excellent agreement with the LQCD data, lying almost entirely within the error bands, in addition to describing a peak of similar height and even converging to the conformal limit at large $\mu_{I}$. The TRS description, on the other hand, is not able to describe the overall behavior because a finite value cannot be assigned to the speed of sound at large isospin chemical potentials. This behavior can be related to the pion condensate, which vanishes at a high $\mu_{I}$ in the TRS, while remaining nonvanishing within the MSS prescription, as shown in Fig. 2. The persistence of the pion condensate up to high values of $\mu_{I}$ in the MSS shown in Fig. 2 might be seen as an extrapolation, where we are going beyond the validity of the effective model. However, even at moderate values of $\mu_{I}$ we can clearly see evident differences between TRS and MSS. As we show next, these subtle differences between the schemes at moderate $\mu_{I}$ are enough to capture the correct behavior for $c_{s}^{2}$ in the MSS case. Nonetheless, Ref. Abbott:2024vhj estimates the pion condensate through a subtraction of the pQCD pressure without a gap from the LQCD pressure, finding that it still contributes significantly in the $\mu_{I}\in[1500,3000]$ MeV region.
@@
[…]
In order to verify the dependence of our findings on the regularization method applied to the vacuum integrals, in Fig. 3 we have also considered the result when using the Pauli-Villars (PV) regularization method within the MSS. We considered only the SU(2) case, which is representative enough as argued in the discussion of the results. For PV, the divergences can be of quadratic or logarithmic order in momentum. Hence, two additional terms are needed in the MSS scheme, more specifically, divergent loop functions are expressed as
$f(k,M_{0})\to f(k,M_{0})+C_{1}f(k,M_{1})+C_{2}f(k,M_{2})\,$, (6)
where $M_{i}^{2}=M_{0}^{2}+\alpha_{i}\Lambda_{\rm PV}^{2}$. We choose $C_{1}=1$, $C_{2}=-2$, $\alpha_{1}=2$ and $\alpha_{2}=1$, as commonly adopted in the literature. The input parameters shown in Table 1 remain the same for PV and the corresponding outputs are: $\Lambda_{\rm PV}=863.3~{\rm MeV}$, $G=3.708~{\rm GeV}^{-2}$ and $m_{l}=5.205~{\rm MeV}$. In Fig. 3 we observe a good description of the GP-model data also for the PV-MSS case, with a slightly smoother curve than in the MSS-sharp cutoff case. Both the qualitative peak structure and the convergence to the conformal limit remain in the PV-MSS as well, thus indicating that the $c_{s}^{2}$ peak is a robust feature of MSS, rather than an artifact of a particular regularization of the vacuum. Furthermore, in all cases within the MSS, the height of the peak respects the upper bound $c_{s}^{2}=0.781$ of nuclear matter given in Ref. Hippert:2024hum , which was obtained through the calculation of hydrodynamic transport coefficients from first principles.
@@
Conclusions. In this paper, we investigated the importance of the regularization scheme for the description of isospin imbalanced matter in the context of the NJL model. Of particular interest is the existence of a peak in the speed of sound and recently observed by lattice QCD simulations Brandt:2022hwy ; Abbott:2023coj ; Abbott:2024vhj . We worked within the MSS, which disentangles medium contributions from the vacuum such that only the latter is regularized. The results were compared with those obtained by the usual TRS procedure, which is typically used in NJL-type models in the literature.
@@
Throughout the analysis, it was made clear that the speed of sound peak is a natural prediction of the MSS. One must keep in mind though that the NJL model has limited applicability at high energy scales as it is nonrenormalizable. Consequently, results with an isospin chemical potential greater than the regularization scale $\Lambda$ must be taken with great care. In fact, the TRS results are problematic at high $\mu_{I}$ and do not recover the conformal limit for the speed of sound. However, remarkably, despite the shortcomings expected from an effective model, the MSS results asymptotically approach $c_{s}^{2}=1/3$ as $\mu_{I}$ increases. This feature may be related to the behavior of the pion condensate, which is nonvanishing in this regime, in contrast to the prediction from the TRS prescription. Although concerned with superconducting systems, the renormalization group consistent regularization proposed in Ref. Gholami:2024diy finds a nonvanishing diquark condensate at large chemical potential, which is not the case with a conventional regularization. In this sense, the matching behaviors seem to signal that the MSS plays an important role in minimizing the dependence of physical quantities on the regularization scale. As the pion condensate is a central quantity in the analysis, a comparison with lattice results from Refs. Brandt:2017oyy ; Brandt:2018wkp would be valuable. The aforementioned simulations, however, were performed at finite temperature, which requires a separate study from what we have done in the present paper  [^3] We note, however, that Ref. Brandt:2018bwq explored the case of temperatures significantly below the relevant QCD scale, in which case the results presented in that reference can be well approximated to $T=0$. Although the pion condensate is not explicitly presented there, relevant information can be extracted from their Fig. 3 and also using their Eq. (A11).. An extension of this analysis that includes the pion condensate and also at finite temperature is being carried out and will be reported in future work.
In summary, the MSS has consistently demonstrated reliability across systems free from the sign problem, showing excellent agreement with lattice QCD where traditional cutoff-based schemes fail. Its successful applications extend to regimes inaccessible to lattice simulations, such as two-color-superconducting quark matter, where it yields results consistent with effective models. The present work advances this framework by showing that MSS naturally accounts for the peak in the speed of sound at finite isospin density, in agreement with recent findings from both lattice QCD and functional renormalization group approaches. For this reason, we believe that a comparison with the RG-consistent method at finite isospin density would be interesting and could potentially help to further validate the applicability of the medium separation scheme also at the high density regime.
@@
Acknowledgments. This work was partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grants No. 312032/2023-4 (R.L.S.F.), No. 141270/2023-3 (B.S.L.) and No. 307286/2021-5 (R.O.R.); Fundação de Amparo à Pesquisa do Estado do Rio Grande do Sul (FAPERGS), Grants No. 24/2551-0001285-0 (R.L.S.F.), No. 23/2551-0000791-6, and No. 23/2551-0001591-9 (D.C.D); Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021 (R.O.R.). The work is also part of the project Instituto Nacional de Ciência e Tecnologia—Física Nuclear e Aplicações (INCT—FNA), Grant No. 464898/2014-5 and supported by the Serrapilheira Institute (Grant No. Serra-2211-42230).
Data Availability. The data are not publicly available. The data are available from the authors upon reasonable request.